# Chemist-X: Large Language Model-empowered Agent for Reaction Condition Recommendation in Chemical Synthesis

### v5

@@ Abstract
Recent AI research plots a promising future of automatic chemical reactions within the chemistry society. This study proposes Chemist-X, a transformative AI agent that automates the reaction condition recommendation (RCR) task in chemical synthesis with retrieval-augmented generation (RAG) technology. To emulate expert chemists’ strategies when solving RCR tasks, Chemist-X utilizes advanced RAG schemes to interrogate online molecular databases and distill critical data from the latest literature database. Further, the agent leverages state-of-the-art computer-aided design (CAD) tools with a large language model (LLM) supervised programming interface. With the ability to utilize updated chemical knowledge and CAD tools, our agent significantly outperforms conventional synthesis AIs confined to the fixed knowledge within its training data. Chemist-X considerably reduces chemists’ workload and allows them to focus on more fundamental and creative problems, thereby bringing closer computational techniques and chemical research and making a remarkable leap toward harnessing AI’s full capabilities in scientific discovery.
@@ 1 Introduction
Recent “artificial intelligence (AI) for chemistry” research aims to free human labor with AI-supervised chemical reaction platforms [1, 2, 3, 4]. These innovative platforms employ machine learning (ML) algorithms and data-driven methods to orchestrate synthetic routes and identify optimal reaction conditions for maximizing yields. In the foreseeable future, chemical robots equipped with integrated AIs will autonomously design and refine chemical reactions, thus allowing human experts to focus on more creative and foundational research. Yet, traditional AI systems fall short of human experts who possess a broad spectrum of chemical knowledge and can continually update their understanding by consulting related literature when dealing with new problems. Conventional AI models, being confined to the knowledge from their training data, exhibit limited generalization ability when encountering unknown reactions. They can provide chemists with preliminary assistance for familiar reactions but cannot realize the ambitious aim of fully automated synthesis of unprecedented products.
@@
Recent technical breakthroughs in retrieval-augmented generative AI (RAG-AI) have shed light on this problem. RAG-AI, as its name suggests, harnesses the strengths of retrieval-based methods and the capacities of generative AIs like large language models (LLMs). RAG enables an AI to retrieve relevant information from a vast and continuously updated knowledge source [5, 6], while LLM equips the AI with close-to-human abilities in information analysis and code writing [7, 8]. This paper develops a sophisticated chemical reaction agent, namely Chemist-X, with RAG-AI technology. As a proof of concept, we focus on RAG for reaction condition recommendation (RCR), which is equally important as retrosynthesis but receives less attention in chemical synthetic research.
LLMs like ChatGPT [9] or GPT-4 [10], thanks to their advanced generalization capabilities and near-human intelligence, demonstrate remarkable proficiency in devising feasible execution plans for RCR problems. The plan proposed by the LLM can be summarized as the following methodical three-phase approach: 1) identify structurally analogous molecules, which may share similar chemical characteristics with the target products, to provide foundational references for further analysis; 2) review existing literature about these analogous molecules to narrow down the chemical space of potential reaction conditions; and 3) utilize Computer-Aided Design (CAD) technology to select a subset of promising conditions from the refined chemical space. This systematic ”search-analyze-recommend” pattern suggested by the LLM well aligns with the methodology a human expert might undertake when addressing the RCR challenge associated with an unfamiliar molecule, thereby providing strong evidence for the practicality of the LLM’s three-phase solution. Fig. 1 illustrates the three-phase solution with more technical details.
Figure 1: The three-phase RCR framework of Chemist-X. a) The agent’s problem-solving procedure presented in the form of a pipeline. All three phases are automatically executed with the control Chemist-X. b) A detailed illustration of operations involved in Chemist-X. The LLM is the kernel of the agent. It is important to note that the figure follows the recent concept of “AI agent”, in which an LLM-empowered AI system can achieve close-to-human performance in a specific task owing to its abilities in i) using additional knowledge as “memory”, ii) leveraging “tools” like pre-defined APIs and auto-generated codes, and iii) taking necessary actions so that the processing can automatically move on to the next step.
@@
Phase One retrieves molecules analogous to the target product by leveraging an LLM-empowered code generation scheme tailored to application programming interfaces (APIs) provided by molecular databases such as PubChem [11] and ChemSpider [12]. These databases typically offer APIs to facilitate information retrieval and provide examples of usage in their tutorial handbooks. However, software programming based on API documents requires a related background in computer science, and it could be a time-consuming task given that each database has its own API protocols and documentation. Phase One makes such API-based information retrieval accessible to researchers regardless of their familiarity with software programming, thereby bridging the gap between chemists and computer scientists.
@@
In pursuit of automatic API programming without human intervention, we harness in-context learning (ICL) [13, 14], an effective few-shot learning method that equips the LLM with necessary API knowledge as background information and augments its code generation capabilities using exemplar codes in prompts. It is important to note that providing the LLM with the entire programming handbook in ICL is counterproductive, as this could overwhelm the LLM, making it struggle to pinpoint the most relevant code segments. To address this problem, we put forth the concept of top match slice (TMS), which refers to the portion of the document bearing the utmost semantic similarity to the problem description provided to the LLM. This similarity is measured by the distance between our problem description and the tested document slice within the semantic embedding space [15]. Our experiments indicate that incorporating the TMS of an API handbook into the ICL prompt significantly enhances the LLM’s performance in retrieving information about similar molecules via APIs. The scheme outperforms ICL with complete documentation or random document slices in multiple aspects, including higher code reliability, better cost-efficiency, and reduced response time.
Phase Two retrieves detailed literature information from online sources and analyzes the data with automatically generated code. Upon the completion of Phase One, we possess a list of structurally similar molecules. To understand how previous papers deal with similar reactions, our agent first searches for chemical reactions related to each of these analogous molecules in online literature platforms like PubMed [16] via web crawlers. With HTML content collected by these crawlers, the agent then distills the raw data to construct a potential reaction condition subspace for the target molecules.
Beyond the creation of the internet crawler, Chemist-X also realized an efficient HTML analyzing scheme that overcomes the limitation of the LLM’s input information length. An online literature platform may employ complex web designs, making its HTML source surpass the maximum token length permissible for LLMs. Chemist-X addresses this obstacle by instructing the LLM to generate Python code tailored to HTML analysis, rather than directly feeding the raw HTML source to the LLM as in prior works [17, 18]. Compared with the conventional scheme, our script-based method demonstrates higher accuracy and efficiency in extracting reaction condition data from extensive HTML sources.
@@
Phase Three makes final recommendations. Once the chemical reaction subspace has been refined, our focus shifts to identifying the optimal set of reaction conditions with top yield expectations. There are some prior studies investigating CAD schemes for such recommendation tasks [19, 20, 21]. Given the intense research interest in RCR for chemical synthesis, there may be more advanced CAD schemes in the future. Chemist-X serves as a universal platform for cutting-edge CAD algorithms, wherein the latest research findings could be integrated in the form of APIs. To fulfill this objective, our agent harnesses LLMs to comprehend the functionality of each chemical CAD algorithm via its documentation and to select appropriate CAD tools through API programming. As the chemistry community continues to update CAD tools, our platform enables the continuous evolution of Phase Three, ensuring it remains at the forefront of technological advancement.
@@
This paper developed an advanced ML model for RCR tasks and packaged it into an API so that the agent can execute the new model for concept proving. Specifically, in our new model, we designed a novel chemical reaction fingerprint and applied it to enhance the conventional ML models. The fingerprint diverges from traditional reaction encoding approaches in two key aspects. First, we apply chemistry-informed molecular graph (CIMG) [22], an innovative molecular descriptor based on graph neural networks (GNNs) [23], as the underlying molecule encoding method to boost the quality of molecular descriptions. Second, we integrate supervised contrastive learning (SCL) [24] in our fingerprint-generation network, thus enhancing the model’s ability to distinguish the unique features of high-yield reactions from others. Further, to better serve the SCL process, we replaced the noise-sensitivity numerical descriptions of reaction yields with coarse categorical yield labels. With the above efforts, our reaction embedding scheme, denoted as the coarse-label SCL (CL-SCL) fingerprint, demonstrates exceptional efficacy in capturing the essential characteristics of high-yield reactions. It constantly outperforms prior studies in terms of precision and robustness across a variety of ML models.
@@
In general, the three-phase framework supervised by LLM, along with the specific techniques employed in each phase, enables Chemist-X to automatically acquire the latest online molecules/literature knowledge and to utilize the most cutting-edge CAD tools. With the advantages of continuous evolution and unmanned processing, Chemist-X makes a significant contribution to the AI-chemistry community and makes a meaningful step towards the ultimate goal of AI-supervised chemical reaction platforms.
@@ 2 Results and Discussions
This section demonstrates how Chemist-X completes the RCR task with an example Suzuki reaction targeting on 6-(1-methyl-1H-indazol-4-yl) quinoline, whose SMILES is written as CN1C2=CC=CC(C3=CC=C4N=CC=CC4=C3)=C2C=N1. We investigate the agent’s performance in each phase via comprehensive unit tests, and then we conduct wet-lab experiments to validate the recommendation results.
2.1 Unit Tests
2.1.1 Unit Tests for Phase One
Phase One employs ICL to craft Python code capable of interfacing with PubChem via its official APIs. Figure 2a illustrates our prompt, which includes a fixed part and a flexible part. Our question is described in the fixed part, while the reference example of ICL should be contained in the flexible part. Our agent identifies the most pertinent section of the API documentation and uses it to fill in the flexible part of the prompt.
@@ 2.1.1 Unit Tests for Phase One
The agent’s accuracy and efficiency are substantially improved by incorporating the TMS. We compare our TMS-ICL approach against three alternatives: (i) zero-shot learning without any background information or examples, (ii) ICL using the complete API documentation as background reference, and (iii) ICL with a randomly chosen slice as background reference. We conduct ten trials of our TMS-ICL and the three alternative methods using a GPT-4 model. We assess the performance of these four code-generation strategies based on computational resources used and the success rate in generating functionally correct code in the ten trials. Specifically, we consider four aspects of computational requirements: the average number of question/answer tokens used in one interaction with the GPT-4 model, the model’s average response time in completing one code design task, and the average cost of utilizing the commercial LLM model for one time. Experimental results presented from Figure 2d to Figure 2h indicate that our TMS-ICL method consumes slightly more computational resources than the random slice approach yet delivers a remarkably higher success rate than all three alternatives. The zero-shot approach uses the least computation resources but has a disappointingly low success rate. The all-document approach shows modest improvement over zero-shot learning but does not match the performance of our TMS-ICL and requires substantially more computational resources.
@@
2.1.2 Unit Tests for Phase Two
Phase Two retrieves reaction-related information from literature platforms and extracts chemical reaction condition data, such as bases and solvents, from HTML sources. We develop an innovative prompting framework for generating web crawlers via ICL (see 3a) and an HTML analysis scheme that employs LLMs to generate Python code tailored to information extraction from HTML contents (see 3b).
We apply GPT-4 in generating the web crawler. We note that GPT-4 offers customized parameter configuration for users’ flexible adaption. Among these parameters, temperature and presence penalty play a vital role in code generation: the former affects the creativeness of the output, while the latter influences the repetition of generated content. In Figure 3a, we investigate how these two parameters affect the model’s performance in our case. As we can see, our task is sensitive to the model’s temperature and performs well when the temperature is low, given that being focused and deterministic is more important than creative thinking in our web crawling task.
As for HTML analysis, experiments presented in Figure 3d demonstrate that our approach significantly outperforms conventional HTML analysis methods that input the entire HTML source into LLMs directly [17, 18]. Our experiments focus on the top 30 structurally similar molecules identified in Phase One. The HTML source of each molecule’s search result has been collected by the crawler. Upon manual review, these files include 143 pieces of reaction condition data in total. The conventional approach requires 30 separate interactions with the GPT-4 model, i.e., one round of interaction for each HTML file; while our method involves a single interaction for code generation, which is then repeatedly executed to extract information from these 30 files. Experimental results prove that our scheme consumes significantly fewer question/answer tokens than the conventional scheme in the 30-HTML task, which further results in fewer costs and shorter response time. Despite consuming fewer computational resources, our method achieves superior extraction accuracy, successfully identifying in total 141 out of 143 pieces of correct data from these HTML sources, while the conventional scheme only figures out 78 pieces of data. Finally, we calculate the F1 score of the two schemes: 70.59% for the conventional scheme and 99.30% for our scheme.
Figure 3: Comparative Analysis of HTML Extraction Methods. a) The prompt we used for developing the web crawler. b) The prompt we used for developing the HTML analyzer. Here we apply the ICL method with three examples and hint the LLM to apply a character-matching strategy when generating the code. c) A 3D figure showing how the generation process of the web crawler is affected by the LLM’s parameter. d) A benchmarking between our code-based HTML analysis method with the conventional method that inputs HTML directly into the LLM. Notably, to make the comparison possible, we simplified the challenge of the conventional scheme by pre-processing the raw HTML to remove extraneous content, such as formatting and search bar. Otherwise, the length of some raw HTML sources would exceed GPT-4’s token limit. Despite this adjustment, our method still demonstrates higher precision and lower resource usage.
2.1.3 Unit Tests for Phase Three

### v6

# Chemist-X: Large Language Model-Powered Agent for Recommending Reaction Conditions in Chemical Synthesis and Autonomous Laboratories

@@ Abstract
Recent AI research plots a promising future of automatic chemical reactions within the chemistry society. This study proposes Chemist-X, a comprehensive AI agent that automates the reaction condition optimization (RCO) task in chemical synthesis with retrieval-augmented generation (RAG) technology and AI-controlled wet-lab experiment executions. To begin with, as an emulation on how chemical experts solve the RCO task, Chemist-X utilizes a novel RAG scheme to interrogate available molecular and literature databases to narrow the searching space for later processing. The agent then leverages a computer-aided design (CAD) tool we have developed through a large language model (LLM) supervised programming interface. With updated chemical knowledge obtained via RAG, as well as the ability in using CAD tools, our agent significantly outperforms conventional RCO AIs confined to the fixed knowledge within its training data. Finally, Chemist-X interacts with the physical world through an automated robotic system, which can validate the suggested chemical reaction condition without human interventions. The control of the robotic system was achieved with a novel algorithm we have developed for the equipment, which relies on LLMs for reliable script generation. Results of our automatic wet-lab experiments, achieved by fully LLM-supervised end-to-end operation with no human in the lope, prove Chemist-X’s ability in self-driving laboratories.
@@ 1 Introduction
Recent “artificial intelligence (AI) for chemistry” research aims to free human labor with AI-supervised chemical reaction platforms [1, 2, 3, 4]. These innovative platforms employ machine learning (ML) algorithms and data-driven methods to orchestrate synthetic routes and identify optimal reaction conditions for maximizing yields. In the foreseeable future, chemical robots equipped with integrated AIs will autonomously design and refine chemical reactions, as well as automatically perform chemical reactions, allowing human experts to be freed from repetitive and time-consuming experimental operations and focus on more creative and foundational research. Yet, traditional AI systems fall short of human experts who possess a broad spectrum of chemical knowledge and can continually update their understanding by consulting related literature when dealing with new problems. Conventional AI models, being confined to the knowledge from their training data, exhibit limited generalization ability when encountering unknown reactions. These AIs can provide chemists with preliminary assistance for familiar reactions but cannot realize the ambitious aim of fully automated synthesis of unprecedented products.
@@
Recent technical breakthroughs in retrieval-augmented generative AI (RAG-AI) have shed light on this problem. RAG-AI, as its name suggests, harnesses the strengths of retrieval-based methods and the capacities of generative AIs like large language models (LLMs). RAG enables an AI to retrieve relevant information from a vast and continuously updated knowledge source [5, 6], while LLM equips the AI with close-to-human abilities in information analysis and code writing [7, 8]. This paper develops a sophisticated chemical reaction agent, namely Chemist-X, with RAG-AI technology. We focus on RAG for reaction condition optimization (RCO), which is equally important as retrosynthesis but receives less attention in synthetic research.
We have designed the following three-phase pipeline to leverage LLMs’ generalization capabilities and near-human intelligence. Fig. 1(a) illustrates the three-phase solution with technical details, which can be summarized as follows. In Phase One, Chemist-X uses the LLM to generate usable python code for searching existing chemical literature and molecule datasets for reaction information about the target molecule (and that of its analogous molecules) to narrow down the chemical space of potential reaction conditions. Then in Phase Two, Chemist-X, as an intelligent agent supported by LLM, utilizes Computer-Aided Design (CAD) tools to select a subset of promising conditions from the refined chemical space obtained in Phase One. Finally in Phase Three, our agent controls auto-lab equipment (via LLM generated control script) to conduct wet-lab experiments to validate potential reaction conditions suggested by Phase Two. This systematic “retrieve-recommend-validate” pattern applied in our system well aligns with the methodology a human expert might undertake when addressing the RCO challenge associated with an unfamiliar molecule, and we will later show that LLMs can well address the executive challenges in each phase owing to their strong task analysis and code generating ability. The combination of the novel three-phase pipeline and the intelligence of LLMs makes the system’s reliable performance in practical tasks become possible.
Figure 1: The three-phase RCO framework of Chemist-X, in which all phases are automatically executed under the control of the LLM agent.
We now introduce the execution of each phase. Phase One applies a hierarchical matching strategy to efficiently retrieve reaction conditions to a target product from both online and offline databases. Our retrieval method could be applied on various public or private databases. For concept-proving purposes, we collaborated with the Chemical Abstracts Service (CAS) organization and built a private research-use chemistry database on the Amazon Elastic Compute Cloud (EC2) platform, in which all the data has been properly licensed with CAS. Hierarchical Matching Strategy begins with an exact SMILES-matching policy to check if the target compound exists in our database. If found, the reaction conditions are retrieved using the Amazon EC2 API [9]. If the compound is not present, the strategy employs the PubChem API [10] to generate a list of similar molecules. The agent searches for the most similar molecule using exact SMILES matching and retrieves its reaction conditions. If unavailable, the process continues with the next similar molecule until successful. If no suitable match is found, the agent moves directly to Phase Two.
@@
Complementing this strategy is an LLM-empowered code generation scheme tailored to APIs such as PubChem and Amazon EC2. These online databases typically offer APIs to facilitate information retrieval and provide examples of usage in their tutorial handbooks. However, software programming based on API documents requires a related background in computer science, and it could be a time-consuming task given that each database has its own API protocols and documentation. Phase One makes such API-based information retrieval accessible to researchers regardless of their familiarity with software programming, thereby bridging the gap between chemists and computer scientists.
@@
In pursuit of automatic API programming without human intervention, we harness in-context learning (ICL) [11, 12], an effective few-shot learning method that equips the LLM with necessary API knowledge as background information and augments its code generation capabilities using exemplar codes in prompts. It is important to note that providing the LLM with the entire programming handbook in ICL is counterproductive, as this could overwhelm the LLM, making it struggle to pinpoint the most relevant code segments. To address this problem, we put forth the concept of top match slice (TMS), which refers to the portion of the document bearing the utmost semantic similarity to the problem description provided to the LLM. This similarity is measured by the distance between our problem description and the tested document slice within the semantic embedding space [13]. Our experiments indicate that incorporating the TMS of an API handbook into the ICL prompt significantly enhances the LLM’s performance in information retrieval. The scheme outperforms ICL with complete documentation or random document slices in multiple aspects, including higher code reliability, better cost-efficiency, and reduced response time.
@@
Phase Two makes the recommendations. Once the chemical reaction subspace has been refined with the effort in phase one, our focus shifts to identifying the optimal set of reaction conditions with top yield expectations. There are some prior studies investigating CAD schemes for such recommendation tasks [14, 15, 16]. Given the intense research interest in RCO for chemical synthesis, there may be more advanced CAD schemes in the future. Chemist-X serves as a universal platform for cutting-edge CAD algorithms, wherein the latest research findings could be integrated in the form of APIs. To fulfill this objective, our agent harnesses LLMs to comprehend the functionality of each chemical CAD algorithm via its documentation and to select appropriate CAD tools through API programming. As the chemistry community continues to update CAD tools, our platform enables the continuous evolution of Phase Two, ensuring it remains at the forefront of technological advancement.
@@
Apart from the general CAD platform provided in Phase Two, this paper developed an advanced ML model for the RCO task. We packaged the algorithm into an API so that the agent can execute the new model for concept proving. Specifically, in our new model, we designed a novel chemical reaction fingerprint and applied it to enhance the conventional ML models. The fingerprint diverges from traditional reaction encoding approaches in two key aspects. First, we apply chemistry-informed molecular graph (CIMG) [17], an innovative molecular descriptor based on graph neural networks (GNNs) [18], as the underlying molecule encoding method to boost the quality of molecular descriptions. Second, we integrate supervised contrastive learning (SCL) [19] in our fingerprint-generation network, thus enhancing the model’s ability to distinguish the unique features of high-yield reactions from others. Further, to better serve the SCL process, we replaced the noise-sensitivity numerical descriptions of reaction yields with coarse categorical yield labels. With the above efforts, our reaction embedding scheme, denoted as the coarse-label SCL (CL-SCL) fingerprint, demonstrates exceptional efficacy in capturing the essential characteristics of high-yield reactions. It constantly outperforms prior studies in terms of precision and robustness across a variety of ML models.
Phase Three introduces a general computer control method that enables our agent to operate wet-lab experimental platforms to validate the reaction conditions recommended in Phase Two. Our motivation is that many wet-lab experimental platforms do not have predefined APIs. Instead, many rely on user interactions (e.g., mouse clicks or touchscreen inputs) within software interfaces for operation. To enhance the generalization capability of our AI agent, we designed a general computer control framework that can simulate human manipulation and learn from human-operated demonstration videos. This is achieved by utilizing the LLM’s reasoning ability, its code generation ability (based on the pyautogui Python package), and the assistance of several tools tailored to the agent: a key frame detection algorithm, a mouse-click detection algorithm, and an optical character recognition scheme.
This paper implements the automation control algorithm developed in Phase Three on a large-scale molecular manufacturing platform known as iChemFoundry (IC). With advantages such as full automation, time efficiency, reduced risk, high reproducibility, and broad versatility, the IC platform effectively addresses the widespread challenges of automating chemical synthesis in laboratory settings [20, 21]. Notably, the IC platform successfully automates reaction pretreatment, including solid dosing and synthesis under inert gas protection—two critical challenges faced by currently reported robotic systems [22].
For testing, we use the Suzuki–Miyaura coupling, a representative chemical reaction in industrial chemistry, as the target reaction. During the execution of these Suzuki–Miyaura reactions, Chemist-X recommends specific reaction conditions, which are then automatically validated on the IC platform, utilizing Unchained Labs, the HPLC characterization island, and integrated robotic arms. The IC platform autonomously executes all experimental operations, including pretreatment with solid and liquid reagents, synthesis under inert gas protection, post-treatment, sample transfer, and reaction characterization, without human intervention. The compelling results from wet-lab experiments demonstrate the feasibility of using LLMs to replicate human interactions with software interfaces for general equipment control.
@@
In general, our three-phase framework supervised by LLM, along with the specific techniques employed in each phase, enables Chemist-X to automatically acquire the latest online molecules/literature knowledge, utilize the most cutting-edge CAD tools, and control automatic wet-lab equipment for reaction validation. With these advantages, Chemist-X makes a significant contribution to the AI-chemistry community and makes a meaningful step towards the ultimate goal of AI-supervised chemical reaction platforms.
@@ 2 Results and Discussions
This section demonstrates how Chemist-X completes the RCO task with an example Suzuki reaction targeting 6-(1-methyl-1H-indazol-4-yl) quinoline, whose SMILES is written as CN1C2=CC=CC(C3=CC=C4N=CC=CC4=C3)=C2C=N1. We now investigate the agent’s performance in Phase One and Phase Two via comprehensive unit tests, and then we validate the generated recommendation results via wet-lab experiments controlled by Phase Three.
2.1 Phase One Unit Tests
Phase One employs ICL to craft Python code capable of interfacing with Amazon EC2 and PubChem via its official APIs. Since the code generation pipeline for the Amazon EC2 platform is highly similar to those of the PubChem API, this subsection uses the PubChem API for demonstration to fulfill the page limit. Figure 2a illustrates the prompt we used for the code generation pipeline, which includes a fixed part and a flexible part. Our question is described in the fixed part, while the reference example of ICL should be contained in the flexible part. Our agent identifies the most pertinent section of the API documentation and uses it to fill in the flexible part of the prompt.
@@ 2.1 Phase One Unit Tests
The agent’s accuracy and efficiency are substantially improved by incorporating the TMS. We compare our TMS-ICL approach against three alternatives: (i) zero-shot learning without any background information or examples, (ii) ICL using the complete API documentation as background reference, and (iii) ICL with a randomly chosen slice as background reference. We conduct ten trials of our TMS-ICL and the three alternative methods using a GPT-4 model. We assess the performance of these four code-generation strategies based on computational resources used and the success rate in generating functionally correct code in the ten trials. Specifically, we consider four aspects of computational requirements: the average number of question/answer tokens used in one interaction with the GPT-4 model, the model’s average response time in completing one code design task, and the average cost of utilizing the commercial LLM model for one time. Experimental results presented from Figure 2d to Figure 2h indicate that our TMS-ICL method consumes slightly more computational resources than the random slice approach yet delivers a remarkably higher success rate than all three alternatives. The zero-shot approach uses the least computational resources but has a disappointingly low success rate. The all-document approach shows modest improvement over zero-shot learning but does not match the performance of our TMS-ICL and requires substantially more computational resources.
@@
2.2 Phase Two Unit Tests